# Stability of $L^{p}$ Dirichlet solvability under small bi-Lipschitz transformations of domains

### v1

@@ 1.1. Motivation and statement of main results
For convex domains, as one has a pointwise bound on the gradient of the Green function up to the boundary, one can obtain $(D)_{p}$ for the full range of $p$ using the Hardy-Littlewood maximal function (this is observed in [She06]). Combining this idea with Schauder theory, we can extend the result slightly beyond convex domains, namely, to domains whose boundary is (locally) the image of a convex graph under a $C^{1,\alpha}$ diffeomorphism (see Appendix B). However, since the argument relies heavily on boundary $C^{1}$ estimates of solutions, reducing the regularity of the $C^{1,\alpha}$ diffeomorphism to only $C^{1}$ is beyond reach of this method.
@@ 1.3. Main ideas of our approach
We conclude the introduction by noting that our approach appears robust enough to apply to more general elliptic operators. For instance, we are hopeful that our results could be extended to operators satisfying the DKP condition, although substantial modifications would be required. Our results may also extend to more general domains whose boundaries are not necessarily given by graphs, following the ideas of [KT03]. In a different direction, our method might be adapted to address Question 1.2 for the Neumann problem; however, due to the inherent difficulty of the Neumann problem, new ideas would be needed. Finally, a precise characterization of the necessary condition in Question 1.1 remains open and is likely to be highly challenging.
@@ 2.1. Notations and definitions
Let $\Omega$ be a Lipschitz domain in $\mathbb{R}^{n}$. We say a quantity $f$ defined on $\Omega$ satisfies the Carleson measure condition with sup (denoted by $f\in\operatorname{CM_{sup}}$) if $\sup_{\frac{1}{2}B_{X}}|f|^{2}\frac{1}{\delta(X)}dX$ is the density of a Carleson measure on $\Omega$, that is, there is some constant $M>0$ such that
$\iint_{B(x,r)\cap\Omega}\left(\sup_{Z\in\frac{1}{2}B_{Y}}|f(Z)|\right)^{2}\frac{dY}{\delta(Y)}\leq Mr^{n-1}\quad\text{for all }x\in\partial\Omega,r>0.$
@@
We say a quantity $f$ defined on $\Omega$ satisfies the Carleson measure condition (denoted by $f\in\operatorname{CM}$) if $|f(X)|^{2}\frac{1}{\delta(X)}dX$ is the density of a Carleson measure on $\Omega$.
@@
Observe that if $f\in\operatorname{CM_{sup}}(M)$, then there exists $C>0$ such that $|f(X)|\leq CM^{1/2}$ for $X\in\Omega$.
@@
is a “cone” with vertex $x$ and aperture $a$.
@@ 2.2. Some properties on $(D)_{p}$
Proof: It is well-known that $(i)$ is equivalent to $\omega_{L}\ll\sigma$ and that there exists $C_{p}>0$ such that the function $k:=\frac{d\omega_{L}}{d\sigma}$ satisfies (2.13), or, in the bounded case, $k^{X_{0}}:=\frac{d\omega_{L}^{X_{0}}}{d\sigma}$ satisfies (2.14). We claim that $k$ and $\kappa$ are comparable, and that $k^{X_{0}}$ and $\kappa^{X_{0}}$ are comparable, with the implicit constants depending only the dimension $n$, the Lipschitz constant $M$ of $\Omega_{0}$, and the ellipticity constant $C_{ellip}$ of $L$. Note that the equivalence follows once this claim is justified.
@@
3. The change of variables - the heart of the matter
@@ 3.1. Set-up estimates on Lipschitz graph domains
Proof: Recall that we write $\partial_{n}G$ for $\left<\nabla G(X),e_{n}\right$>. We can assume without loss of generality that $g(0)=0$, that is, $0\in\partial\Omega_{0}$.
$G^{Y}(X):=\int_{\Omega_{0}}\eta_{\delta(Y)}(Y-Z)G(X,Z)\,dZ.$
[…]
where $K^{*}:=\bigcup_{Z\in K}\overline{B(Z,\delta(Z)/2)}$, and the final convergence is by (3.4).
We claim that
[…]
Integration by parts in $Z_{n}$ gives that
Consequently, $\partial_{n}G^{Y}+H^{Y}-u_{\partial_{n}G^{Y}}$ is a function with non-tangential bounds and zero trace, and is a solution to $\Delta u=0$ in $\Omega_{0}$ by (3.6). Uniqueness of the Dirichlet problem (see [Ken94, Theorem 1.4.4], or [HLMP22, Section 6] for unbounded domains) gives the formula:
$\partial_{n}G^{Y}=u_{\partial_{n}G^{Y}}-H^{Y}\quad\text{in }\Omega_{0}.$
$\partial_{n}G^{Y}(x)=\lim_{h\to 0\atop h>0}\frac{G^{Y}(x+he_{n})}{h}\geq 0.$
Second, we compare $H^{Y}$ and $G^{Y}$ outside of $B_{Y}$. Observe that
by integration by parts. Therefore,
$|H^{Y}(X)|\leq\int_{B_{Y}}\eta_{\delta(Y)}(Y-Z)|\nabla_{Z}G(X,Z)|dZ.$
If $B_{X}$ and $B_{Y}$ do not intersect, we apply Lemma 2.19 to $G(X,\cdot)$ in $\Omega_{0}\setminus B_{X}$. For $Z\in B_{Y}$, $\operatorname{dist}(Z,\partial(\Omega_{0}\setminus B_{X}))=\min\left\{\delta(Z),\operatorname{dist}(Z,B_{X})\right\}$. Obviously, $\delta(Z)\geq\delta(Y)/2$. We check that $\operatorname{dist}(Z,B_{X})\geq\max\left\{\delta(X)/2,|Y-X|-\delta(X)/2-\delta(Y)/2\right\}$. If $\delta(X)<\delta(Y)/4$, then $|Y-X|\geq\delta(Y)-\delta(X)$, and so $|Y-X|-\delta(X)/2-\delta(Y)/2\geq\delta(Y)/8$. This implies that $\operatorname{dist}(Z,B_{X})\geq\delta(Y)/8$, and thus for $Z\in B_{Y}$, $\operatorname{dist}(Z,\partial(\Omega_{0}\setminus B_{X}))\geq\delta(Y)/8$. So we have that
(3.7) $|H^{Y_{i}}(X)|\lesssim 2^{-i}G^{Y_{i}}(X)\lesssim cG^{Y_{i}}.$
For $c>0$, using successively Step 1a, Step 1b, the positivity of $u_{\partial_{n}G^{Y_{i}}}$ and (3.7) we have then that
Since the bound is for all $c>0$, we conclude that
$\frac{\partial_{n}G(X)}{G(X)}\geq 0$
as desired.
(3.8) $\partial_{n}G>0.$
(3.9) $\partial_{n}G(X)\geq cG(X)/\delta(X)\qquad\text{ for all }X\in\Omega_{0}.$
This claim easily yields (H1)–(H2).
To show (3.9), it suffices to show that there is $c>0$ depending only on $n$ and the Lipschitz constant of $g$, such that $\partial_{n}G(0,1)>cG(0,1)$ (recall that $0\in\partial\Omega_{0}$). We can assume without loss of generality that $G(0,1)=1$. By the Harnack inequality, $0<G(X)\lesssim 1$ for all $X\in B(0,2)\cap\Omega_{0}$, with the implicit constant depending only on the dimension and the Lipschitz constant. Moreover, since $G$ is smooth (and in particular, continuous) in $B(0,1)\cap\Omega_{0}$, there exists $c_{0}>0$ (depending only on $n$ and the Lipschitz constant of $g$) such that $G(0,c_{0})\leq 1/2$. So we have that
$\frac{1}{2}\leq G(0,1)-G(0,c_{0})=\int_{c_{0}}^{1}\partial_{n}G(0,s)ds.$
By the mean value theorem, there exists $s_{0}\in(c_{0},1)$ such that $\partial_{n}G(0,s_{0})\geq 1/2$. Finally, since $\partial_{n}G$ is harmonic and nonnegative, the Harnack inequality implies that $\partial_{n}G(0,1)\geq c$ for some $c>0$. $\square$
@@
On top of that, we set $\Omega:=\Phi(\Omega_{0})$, where $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a bi-Lipschitz map satisfying $\|\nabla\Phi-I\|_{\infty}\leq\epsilon$, where $\epsilon>0$ will be small. We want to construct a “better” bi-Lipschitz change of variable $\rho$ from $\Omega_{0}$ to $\Omega$, which preserves the structure of $\Omega_{0}$. By preserving the structure of $\Omega_{0}$, we mean that at least the Jacobian $\nabla\rho$ is close to identity, in the sense that the errors are Carleson measures on $\Omega_{0}$ with small constant.
We view $\Omega$ as a perturbation of $\Omega_{0}$, and the quantity that we want to control is $\lambda:\Omega_{0}\to\mathbb{R}^{n}$, which is a smooth version of the perturbation defined by
where
(3.17) $h(z):=\Phi(z,g(z))-(z,g(z)),$
and
(3.18) $t(X):=\frac{G(X)}{|\nabla G(X)|},\qquad y(X):=\pi\Big(X-\frac{G(X)(\nabla G(X))^{T}}{|\nabla G(X)|^{2}}\Big)=\pi\Big(X-t(X)v_{n}(X)\Big),$
and $v_{n}:=\frac{(\nabla G)^{T}}{|\nabla G|}$ as in Lemma 3.11. Recall that $\pi$ is the orthogonal projection on the first (n1)(n-1)-coordinates.
We emphasize that in our definition, $t:\Omega_{0}\to\mathbb{R}$ is a function and should not be confused with the variable $t\in\mathbb{R}$. We deliberately choose this notation, as we wish to view the function as playing the same role, in $\Omega_{0}$, of the special variable $t$ in $\mathbb{R}^{n}_{+}$. In fact, by Lemma 2.19, (H1), Theorem 3.10, and Corollary 3.15, we have that
@@
(3.27) $|I_{11}|=|I_{11}(X)|\lesssim\beta\left(y(X),t(X)\right),$
where
$\beta(y,t):=\frac{1}{t}\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\fint_{B(y,t)}|h(z)-a-zB|\,dz.$
Observe that the function $(y,t)\mapsto\beta(y,t)$ enjoys the following property: for $(y,t)$ and $(y^{\prime},t^{\prime})\in\mathbb{R}^{n-1}\times(0,\infty)$ satisfying the inclusion of balls $B(y^{\prime},t^{\prime})\subset B(y,Kt)$ for some $K\geq 1$ and $t^{\prime}\geq ct$ for some $c>0$, we have $\beta(y^{\prime},t^{\prime})\leq C_{n,K,c}\beta(y,Kt)$. Moreover, since $h$ is a $C\epsilon$-Lipschitz function, Dorronsoro’s Theorem ([Dor85]) gives that $\beta^{2}(y,t)dy\,dt/t$ is a Carleson measure on $\mathbb{R}^{n-1}\times(0,\infty)$, with Carleson norm smaller than $C\epsilon$, that is,
(3.28) $\sup_{(x,r)\in\mathbb{R}^{n-1}\times(0,\infty)}r^{1-n}\int_{B_{\mathbb{R}^{n-1}}(x,r)}\int_{0}^{r}\beta(y,t)^{2}\frac{dt}{t}dy\leq C\epsilon.$
We want to show that (3.28) implies that $|I_{1}|\in\operatorname{CM_{sup}}(C_{M}\epsilon)$. To do that, consider that map $\Upsilon:(y,t)\mapsto(y,t+g(y))$. It is easy to see that $\Upsilon$ is a bi-Lipschitz map from $\mathbb{R}^{n-1}\times(0,\infty)$ to $\Omega_{0}$. For an arbitrary $X\in\Omega_{0}$, we reserve the notations $y=y(X)$, $t=t(X)$. We denote
$X^{\prime}=\Upsilon(y,t)=(y,t+g(y))\in\Omega_{0},\quad y^{\prime}=y(X^{\prime}),\text{ and }t^{\prime}=t(X^{\prime}).$
We claim that the point $X^{\prime}$ is close to $X$ in the sense that the coefficients $\beta(y^{\prime},t^{\prime})$ can be controlled by a multiple of $\beta(y,Kt)$ for some $K\geq 1$. In fact, since $g$ is $M$-Lipschitz, $\delta(X^{\prime})\approx t$, which implies that $t\approx t^{\prime}$ because $\delta(X^{\prime})\approx t^{\prime}$ and $\delta(X)\approx t$ by (3.19). Next, note that $\pi(X^{\prime})=y$, and so $y^{\prime}=y-t^{\prime}\pi(v_{n}(X^{\prime}))$. This implies that $|y^{\prime}-y|=|t^{\prime}\pi(v_{n}(X^{\prime}))|\leq t^{\prime}\approx t$, and hence $B(y^{\prime},t^{\prime})\subset B(y,Kt)$ for some $K\geq 1$. We deduce that $\beta(y^{\prime},t^{\prime})\lesssim\beta(y,Kt)$ from the property of $\beta$.
We now consider an arbitrary point $Z\in W_{X^{\prime}}$, where $W_{X^{\prime}}=B(X^{\prime},\delta(X^{\prime})/4)$ is a Whitney ball in $\Omega_{0}$ centered at $X^{\prime}=\Upsilon(y,t)$. Obviously, $t(Z)\approx\delta(Z)\approx\delta(X^{\prime})\approx t^{\prime}$, and
$|y(Z)-y^{\prime}|\leq|Z-X^{\prime}|+|t(Z)v_{n}(Z)-t^{\prime}v_{n}(X^{\prime})|\lesssim t^{\prime}.$
This again implies that $B(y(Z),t(Z))\subset B(y^{\prime},Kt^{\prime})$ for some $K\geq 1$, and that $\beta(y(Z),t(Z))\lesssim\beta(y^{\prime},Kt^{\prime})$. Therefore, by (3.27) and the above estimates on $\beta$, we have that for any $x\in\mathbb{R}^{n-1}$ and $r>0$,
$\int_{B(x,r)}\int_{0}^{r}\sup_{Z\in W_{\Phi(y,t)}}|I_{11}(Z)|^{2}\,\frac{dt}{t}\,dy\lesssim\int_{B(x,r)}\int_{0}^{r}\sup_{Z\in W_{\Phi(y,t)}}|\beta(y(Z),t(Z))|^{2}\,\frac{dt}{t}\,dy\\ \lesssim\int_{B(x,r)}\int_{0}^{r}|\beta(y^{\prime},Kt^{\prime})|^{2}\,\frac{dt}{t}\,dy\lesssim\int_{B(x,r)}\int_{0}^{r}|\beta(y,K^{2}t)|^{2}\,\frac{dt}{t}\,dy\lesssim\epsilon\,r^{n-1},$
where we have used (3.28) in the last inequality. Using again the fact that $\Upsilon$ is bi-Lipschitz, we obtain that $|I_{1}|\in\operatorname{CM_{sup}}(C_{M}\epsilon)$ as desired. Step 2 follows.
@@
$+\left(\int_{\mathbb{R}^{n-1}}\big(I_{n\times(n-1)}-(v_{n})^{T}\pi(v_{n})\big)\nabla_{y}\eta_{t}\big(y-z\big)h(z)\,dz\right)=:J_{1}+J_{2}+J_{3}.$
We write $\mathbf{C}:=J_{1}+J_{2}$, and $\mathbf{B}:=J_{3}$. The control on $\mathbf{C}$ is like in Step 2: specifically, $J_{1}$ can be treated like $I_{1}$, $J_{2}$ is treated like $I_{2}$ (using, instead of $v_{n}(\nabla y)\in\operatorname{CM_{sup}}(CM)$ like for $I_{2}$, the Carleson estimate (3.24)).
@@
$\int_{\mathbb{R}^{n-1}}(a+zB)\nabla_{y}\partial_{t}\eta_{t}(y-z)\,dz=\int_{\mathbb{R}^{n-1}}(a+zB)\nabla^{2}_{y}\eta_{t}(y-z)\,dz=0,$
@@
$|K_{11}|\leq\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\left|\int_{\mathbb{R}^{n-1}}\nabla_{y}\partial_{t}\eta_{t}\big(y-z\big)\big[h(z)-(a+zB)\big]\,dz\right|\\ \lesssim\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\frac{1}{t^{2}}\fint_{B(y,t)}|h(z)-(a+zB)|\,dz=\frac{\beta(y,t)}{t}$

### v2

# Stability of $L^{p}$ Dirichlet problem under small bi-Lipschitz transformations of domains

@@ 1.1. Motivation and statement of main results
For convex domains, as one has a pointwise bound on the gradient of the Green function up to the boundary, one can obtain $(D)_{p}$ for the full range of $p$ using the Hardy-Littlewood maximal function (this is observed in [She06]). Combining this idea with Schauder theory, we can extend the result slightly beyond convex domains, namely, to domains whose boundary is (locally) the image of a convex graph under a $C^{1,\alpha}$ diffeomorphism (see Appendix C). However, since the argument relies heavily on boundary $C^{1}$ estimates of solutions, reducing the regularity of the $C^{1,\alpha}$ diffeomorphism to only $C^{1}$ is beyond reach of this method.
@@ 1.3. Main ideas of our approach
We conclude the introduction by noting that our approach appears robust enough to apply to more general elliptic operators. For instance, we are hopeful that our results could be extended to operators satisfying the DKP condition, although substantial modifications would be required. Our results may also extend to more general domains whose boundaries are not necessarily given by graphs, following the ideas of [KT97]. In a different direction, our method might be adapted to address Question 1.2 for the Neumann problem; however, due to the inherent difficulty of the Neumann problem, new ideas would be needed. Finally, a precise characterization of the necessary condition in Question 1.1 remains open and is likely to be highly challenging.
@@ 2.1. Notations and definitions
Let $\Omega_{0}$ be a Lipschitz domain in $\mathbb{R}^{n}$. We say a quantity $f$ defined on $\Omega_{0}$ satisfies the Carleson measure condition with sup (denoted by $f\in\operatorname{CM_{sup}}$) if $\sup_{\frac{1}{2}B_{X}}|f|^{2}\frac{1}{\delta(X)}dX$ is the density of a Carleson measure on $\Omega_{0}$, that is, there is some constant $M>0$ such that
$\iint_{B(x,r)\cap\Omega_{0}}\left(\sup_{Z\in\frac{1}{2}B_{Y}}|f(Z)|\right)^{2}\frac{dY}{\delta(Y)}\leq Mr^{n-1}\quad\text{for all }x\in\partial\Omega,r>0.$
@@
We say a quantity $f$ defined on $\Omega_{0}$ satisfies the Carleson measure condition (denoted by $f\in\operatorname{CM}$) if $|f(X)|^{2}\frac{1}{\delta(X)}dX$ is the density of a Carleson measure on $\Omega_{0}$.
@@
Observe that if $f\in\operatorname{CM_{sup}}(M)$, then there exists $C>0$ such that $\|f\|_{L^{\infty}(\Omega_{0})}\leq CM^{1/2}$ for $X\in\Omega_{0}$.
@@
is a “cone” with vertex $x$ and aperture $a$. Note that the constant $C$ depends on the aperture $a$ of the cone. However, the $L^{p}$ solvability of the Dirichlet problem is independent of the aperture $a$, since real variable techniques shows that $\|N_{a}(.)\|_{L^{p}(\partial\Omega)}$ and $\|N_{b}(.)\|_{L^{p}(\partial\Omega)}$ are equivalent norms (see [Ste93] in the Euclidean case, or Lemma 2.1 in [MPT22] for the proof in our setting).
@@ 2.2. Some properties on $(D)_{p}$
Proof: It is well-known that $(i)$ is equivalent to $\omega_{L}\ll\sigma$ and that there exists $C_{p}>0$ such that the function $k:=\frac{d\omega_{L}}{d\sigma}$ satisfies (2.13), or, in the bounded case, $k^{X_{0}}:=\frac{d\omega_{L}^{X_{0}}}{d\sigma}$ satisfies (2.14)—see, for instance, [Ken94, Theorem 1.7.3]. We claim that $k$ and $\kappa$ are comparable, and that $k^{X_{0}}$ and $\kappa^{X_{0}}$ are comparable, with the implicit constants depending only the dimension $n$, the Lipschitz constant $M$ of $\Omega_{0}$, and the ellipticity constant $C_{ellip}$ of $L$. Note that the equivalence follows once this claim is justified.
@@
3. The change of variables—the heart of the matter
@@ 3.1. Set-up estimates on Lipschitz graph domains
This lemma follows from [CS05, Theorem 11.10] and Lemma 2.19. In fact, we can assume without loss of generality that $g(0)=0$. Given $X_{0}=(x_{0},t_{0})\in\Omega_{0}$, by translation we can assume that $X_{0}=(0,t_{0})$. Let $\delta_{0}=\delta_{0}(n,M)>0$ be the constant found in [CS05, Theorem 11.10], where $M$ is the Lipschitz constant of $g$. If $t_{0}<4\delta_{0}M$, then applying [CS05, Theorem 11.10] [^2] There is a typo in the statement, namely, (11.14) should be $u_{x_{n}}(x)\sim u(x)/\operatorname{dist}(x,\partial\Omega)$. directly to $G$ gives
(3.3) $\partial_{n}G(X_{0})\approx\frac{G(X_{0})}{\delta(X_{0})}.$
If $t_{0}>4\delta_{0}M$, one gets the same estimate by applying the theorem to the harmonic function $u(X):=G\left(\frac{t_{0}}{2\delta_{0}M}X\right)$ and taking $X_{0}=(0,2\delta_{0}M)$. Then (H1) follows from (3.3) immediately, while (H2) follows from (3.3) and Lemma 2.19—recall that we write $\partial_{n}G$ for $\left<\nabla G(X),e_{n}\right$>.
For the sake of completeness, we provide an alternative proof of Lemma 3.2 in Appendix B.
@@
The following notation will be used for the rest of the article. First, $v_{n}:=\frac{(\nabla G)^{T}}{|\nabla G|}$ as in Lemma 3.5. Then, we define
(3.10) $t(X):=\frac{G(X)}{|\nabla G(X)|},\qquad y(X):=\pi\Big(X-\frac{G(X)(\nabla G(X))^{T}}{|\nabla G(X)|^{2}}\Big)=\pi\Big(X-t(X)v_{n}(X)\Big).$
where we recall that $\pi$ is the orthogonal projection on the first (n1)(n-1)-coordinates. We emphasize that in our definition, $t:\Omega_{0}\to\mathbb{R}$ is a function and should not be confused with the variable $t\in\mathbb{R}$. We deliberately choose this notation, as we wish to view the function as playing the same role, in $\Omega_{0}$, of the special variable $t$ in $\mathbb{R}^{n}_{+}$. In fact, by Lemma 2.19, (H1), Theorem 3.4, and Corollary 3.9, we have that
@@
Lemma 3.14.
Let $\varphi:\,\mathbb{R}^{n-1}\to\mathbb{R}$ be a Lipschitz function. Define $\beta_{\varphi}(x,r)$ as the $\beta$-number
$\beta_{\varphi}(x,r):=\frac{1}{r}\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\fint_{B(x,r)}|\varphi(z)-a-(z-x)B|\,dz\qquad\text{ for }x\in\mathbb{R}^{n-1}\text{ and }r>0.$
Then
$X\mapsto\beta_{\varphi}(y(X),t(X))\in CM_{sup}(C_{M}\|\nabla\varphi\|_{\infty}),$
where $t(X)$ and $y(X)$ are as in (3.10), and where $C>0$ is independent of $\varphi$ (depends only on the dimension $n$, and the Lipschitz constant $M$ of $g$).
Proof: We start from the Dorronsoro theorem ([Dor85]), which states that $|\beta_{\varphi}(y,t)|^{2}dy\,dt/t$ is a Carleson measure on $\mathbb{R}^{n-1}\times(0,\infty)$, with Carleson norm smaller than $C\|\nabla\varphi\|_{L^{\infty}}$, that is,
(3.15) $\int_{B_{\mathbb{R}^{n-1}}(x,r)}\int_{0}^{r}|\beta_{\varphi}(z,s)|^{2}\frac{ds}{s}dz\leq C_{Dor}\|\nabla\varphi\|_{\infty}r^{n-1}\qquad\text{ for $x\in\mathbb{R}^{n-1}$ and $r>0$.}$
We want to prove that
(3.16) $\int_{B(x,r)\cap\Omega_{0}}|\beta_{\varphi}(y(X),t(X))|^{2}\frac{dX}{\delta(X)}\leq C\|\nabla\varphi\|_{\infty}r^{n-1}\qquad\text{ for $x\in\partial\Omega_{0}$ and $r>0$.}$
The argument is straightforward and relies on Fubini’s theorem. The details are given below.
Take $X\in\Omega_{0}$, $s\in[t(X),2t(X)]$, and $z\in B_{\mathbb{R}^{n-1}}(\pi(X),t(X))$. From the expression of $y(X)$ in (3.10), one can see that
$|y(X)-\pi(X)|\leq\pi(t(X)v_{n}(X))\leq t(X),$
hence $|z-y(X)|\leq 2t(X)\leq 2s$, and then that
(3.17) $B(y(X),t(X))\subset B(z,3s).$
Observe that the function $(y,t)\mapsto\beta_{\varphi}(y,t)$ enjoys the following property: for $(y,t)$ and $(y^{\prime},t^{\prime})\in\mathbb{R}^{n-1}\times(0,\infty)$ satisfying the inclusion of balls $B(y^{\prime},t^{\prime})\subset B(y,t)$ and $t^{\prime}\geq ct$ for some $c>0$, we have $\beta_{\varphi}(y^{\prime},t^{\prime})\leq C_{n,c}\beta_{\varphi}(y,t)$. So we deduce from (3.17) that
$\beta_{\varphi}(y(X),t(X))\lesssim\beta_{\varphi}(z,3s)\quad\text{ for $z\in B_{\mathbb{R}^{n-1}}(\pi(X),t(X))$ and $s\in[t(X),2t(X)]$},$
and so, for all $X\in\Omega_{0}$
$\beta_{\varphi}((y(X),t(X))\lesssim t(X)^{1-n}\int_{B(\pi(X),t(X))}\int_{t(X)}^{2t(X)}\beta_{\varphi}(z,3s)\,\frac{ds}{s}\,dz\\ \lesssim\left(t(X)^{1-n}\int_{B(\pi(X),t(X))}\int_{3t(X)}^{6t(X)}|\beta_{\varphi}(z,s)|^{2}\,\frac{ds}{s}\,dz\right)^{\frac{1}{2}}.$
Take now $x\in\partial\Omega_{0}$ and $r>0$. By Fubini’s theorem,
$\int_{B(x,r)\cap\Omega_{0}}|\beta_{\varphi}(y(X),t(X))|^{2}\frac{dX}{\delta(X)}\\ \lesssim\int_{B(x,r)\cap\Omega_{0}}t(X)^{1-n}\left(\int_{B(\pi(X),t(X))}\int_{3t(X)}^{6t(X)}|\beta_{\varphi}(z,s)|^{2}\,\frac{ds}{s}\,dz\right)\frac{dX}{\delta(X)}\\ =\int_{\mathbb{R}^{n-1}}\int_{0}^{\infty}|\beta_{\varphi}(z,s)|^{2}\frac{ds}{s}\,dz\Big(\underbrace{\int_{B(x,r)\cap\Omega_{0}}\frac{t(X)^{1-n}}{\delta(X)}{\mathds{1}}_{B(\pi(X),t(X))}(z){\mathds{1}}_{[3t(X),6t(X)]}(s)\,dX}_{=:f(z,s)}\Big).$
We claim that
(3.18) $f(z,s)\lesssim{\mathds{1}}_{B(\pi(x),C_{M}r)}(z){\mathds{1}}_{(0,C_{M}r)}(s).$
for some $C_{M}\geq 2$ independent of $\varphi$, $z$ and $s$. Once the claim (3.18) is established, we deduce that
$\int_{B(x,r)\cap\Omega_{0}}|\beta_{\varphi}(y(X),t(X))|^{2}\frac{dX}{\delta(X)}\lesssim\int_{B(\pi(x),C_{M}r)}\int_{0}^{C_{M}r}|\beta_{\varphi}(z,s)|^{2}\frac{ds}{s}\,dz\\ \leq C_{Dor}\|\nabla\varphi\|_{\infty}(C_{M}r)^{n-1}\lesssim\|\nabla\varphi\|_{\infty}r^{n-1},$
meaning that (3.16) holds as desired for the lemma.
It remains to prove (3.18). First, we observe that $f(z,s)\neq 0$ only if there exists $X\in B(x,r)\cap\Omega_{0}$ such that $|z-\pi(X)|\leq t(X)$ and $s\in[3t(X),6t(X)]$. By (3.11), we have then
$s\leq 6t(X)\lesssim\delta(X)\leq r$
and
$|z-\pi(x)|\leq|z-\pi(X)|+|\pi(X-x)|\leq t(X)+r\lesssim r,$
meaning that there indeed exists $C_{M}\geq 2$ such that $f(z,s)\equiv 0$ outside of $B(\pi(x),C_{M}r)\times(0,C_{M}r)$.
Note also that by (3.11) and the fact that $t(X)\approx s$,
$f(z,s)\lesssim s^{-n}|\{X\in\Omega_{0},\,z\in B(\pi(X),t(X)),\,s\in[3t(X),6t(X)]\}|.$
However, $z\in B(\pi(X),t(X))$ and $s\in[3t(X),6t(X)]$ implies that
(3.19) $|\pi(X)-z|\leq t(X)\leq\frac{s}{3}$
and
$|X-(\pi(X),g(z))|\leq|X-(\pi(X),g(\pi(X)))|+|g(z)-g(\pi(X))|\leq M\delta(X)+M|\pi(X)-z|$
since $g$ is $M$-Lipschitz. By invoking (3.11) again, the latter bound can be continued as follows
$|X-(\pi(X),g(z))|\leq M\delta(X)+M\frac{s}{3}\lesssim s.$
This together with (3.19) gives that
$\left|X-(z,g(z))\right|\leq|\pi(X)-z|+|X-(\pi(X),g(z))|\lesssim s.$
It means that there exists $C$ such that
$\{X\in\Omega_{0},\,z\in B(\pi(X),t(X)),\,s\in[3t(X),6t(X)]\}\subset B((z,g(z)),Cs),$
and thus
$f(z,s)\lesssim s^{-n}|B((z,g(z)),Cs)|\lesssim 1.$
The claim (3.18) and the lemma follows. $\square$
We set now $\Omega:=\Phi(\Omega_{0})$, where $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a bi-Lipschitz map satisfying $\|\nabla\Phi-I\|_{\infty}\leq\epsilon$, where $\epsilon>0$ will be small. We want to construct a “better” bi-Lipschitz change of variable $\rho$ from $\Omega_{0}$ to $\Omega$, which preserves the structure of $\Omega_{0}$. By preserving the structure of $\Omega_{0}$, we mean that at least the Jacobian $\nabla\rho$ is close to identity, in the sense that the errors are Carleson measures on $\Omega_{0}$ with small constant.
We view $\Omega$ as a perturbation of $\Omega_{0}$, and the quantity that we want to control is $\lambda:\Omega_{0}\to\mathbb{R}^{n}$ defined by
where $h$ is the displacement
(3.21) $h(z):=\Phi(z,g(z))-(z,g(z)).$
@@
(3.27) $|I_{11}|=|I_{11}(X)|\lesssim\beta_{h}\left(y(X),t(X)\right)\in CM_{sup}(C_{M}\epsilon),$
by Lemma 3.14 and (3.22). Step 2 follows.
@@
$+\left(\int_{\mathbb{R}^{n-1}}\big(I_{n\times(n-1)}-(v_{n})^{T}\pi(v_{n})\big)\nabla_{y}\eta_{t}\big(y-z\big)h(z)\,dz\right)=:J_{1}+J_{2}+{\bf B},$
where
${\bf B}:=\left(\int_{\mathbb{R}^{n-1}}\big(I_{n\times(n-1)}-(v_{n})^{T}\pi(v_{n})\big)\nabla_{y}\eta_{t}\big(y-z\big)h(z)\,dz\right).$
We write $\mathbf{C}:=J_{1}+J_{2}$. The control on $\mathbf{C}$ is like in Step 2: specifically, $J_{1}$ can be treated like $I_{1}$, $J_{2}$ is treated like $I_{2}$ (using, instead of $v_{n}(\nabla y)\in\operatorname{CM_{sup}}(CM)$ like for $I_{2}$, the Carleson estimate (3.24)).
@@
$\int_{\mathbb{R}^{n-1}}[a+(z-x)B]\nabla_{y}\partial_{t}\eta_{t}(y-z)\,dz=\int_{\mathbb{R}^{n-1}}[a+(z-x)B]\nabla^{2}_{y}\eta_{t}(y-z)\,dz=0,$
@@
$|K_{11}|\leq\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\left|\int_{\mathbb{R}^{n-1}}\nabla_{y}\partial_{t}\eta_{t}\big(y-z\big)\big[h(z)-a-(z-x)B\big]\,dz\right|\\ \lesssim\inf_{a\in\mathbb{R}^{n},B\in\mathbb{R}^{(n-1)\times n}}\frac{1}{t^{2}}\fint_{B(y,t)}|h(z)-a-(z-x)B|\,dz=\frac{\beta_{h}(y,t)}{t}$